OpenStudy (anonymous):

Sam is observing the velocity of a car at different times. After two hours, the velocity of the car is 54 km/h. After four hours, the velocity of the car is 58 km/h. Part A: Write an equation in two variables in the standard form that can be used to describe the velocity of the car at different times. Show your work and define the variables used. (5 points) Part B: How can you graph the equation obtained in Part A for the first six hours? (5 points) @perl

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

@bohotness

OpenStudy (bohotness):

yes :)

OpenStudy (anonymous):

im stuck on this question if you can help

OpenStudy (bohotness):

i can help you :)

OpenStudy (anonymous):

The question is up there

OpenStudy (bohotness):

say velocity = v, and time = t

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

okay but it says in standard form

OpenStudy (anonymous):

nope

OpenStudy (bohotness):

start by finding the slope

OpenStudy (anonymous):

so y2-y1/x2-x1?

OpenStudy (anonymous):

slope is 1/2

OpenStudy (anonymous):

2/1

OpenStudy (bohotness):

okay but its really (2,54) (4,58) 58-54=4/2= 2

OpenStudy (bohotness):

so the required equation will be of form : v=2t+b you need to find b by plugging in one of the given points

OpenStudy (anonymous):

like what

OpenStudy (bohotness):

v=2t+50

OpenStudy (anonymous):

were did you get 50?

OpenStudy (bohotness):

form your slops

OpenStudy (anonymous):

so now we have to convert to standard form

OpenStudy (bohotness):

\(\huge\color{lime}yes\)

OpenStudy (anonymous):

okay

OpenStudy (bohotness):

okay do you kind of get it ?